# Hybrid Action Based Reinforcement Learning for Multi-Objective Compatible Autonomous Driving

### v2

@@ III-C Uncertainty Estimation and Exploration Strategy
$\begin{split}\sigma_{e,i}^{2}\left({s,o,{a_{o}}}\right)={\rm{Va}}{{\rm{r}}_{j\in[1,\cdots,M]}}\left[{{Q_{ij}}\left({s,o,{a_{o}}}\right)}\right]\left|{{}_{\forall o\in{\cal O}}}\right..\end{split}$ (16)
@@
$\begin{split}\sigma_{e}^{2}\left({s,o,{a_{o}}}\right)=\sum\limits_{i=1}^{N}{{\omega_{i}}\sigma_{e,i}^{2}\left({s,o,{a_{o}}}\right)}\left|{{}_{\forall o\in{\cal O}}}\right..\end{split}$ (17)
@@
$\begin{split}{\cal G}={\nabla_{{a_{o}}}}\sigma_{e}^{2}\left({s,o,{a_{o}}}\right)\left|{{}_{\forall{a_{o}}\sim\mu(s)}}\right..\end{split}$ (18)
@@
$\begin{split}{a_{o}}=\arg{\max_{{a_{o}}\in{{\cal A}^{-}}}}\sigma_{e}^{2}(s,o,{a_{o}})\left|{{}_{\forall o\in{\cal O}}}\right.,\end{split}$ (21)
@@
Similarly, the most exploratory discrete action is the one that maximizes epistemic uncertainty: $\arg{\max_{o\in{\cal O}}}\sigma_{e}^{2}(s,o,{a_{o}})$. However, when the epistemic uncertainty of all discrete actions in the set $\cal O$ is low, relying on epistemic uncertainty to choose actions contributes little to strategy exploration, since the agent is already confident about all actions. Thus, we define an uncertainty threshold $\sigma_{e,th}^{2}$ to ensure the agent adopts a greedy strategy and maximizes reward when its uncertainty is low. Additionally, since the parameters of the critic-networks are randomly initialized and their outputs may fluctuate, the estimation of epistemic uncertainty has fluctuations. We use a probabilistic approach rather than directly selecting the action with maximum uncertainty. Specifically, similar to the Softmax function, the probability of selecting a discrete action is based on its uncertainty value, with the total probability across all actions summing to 1. Therefore, the selection of discrete actions follows the function ${\cal F}$, where $o\sim{\cal F}(s,o,{a_{o}})\left|{{}_{\forall o\in{\cal O}}}\right.$:
@@
$\begin{split}\varepsilon\left({s,o,{a_{o}}}\right)=\frac{{{e^{\sigma_{e}^{2}(s,o,{a_{o}})}}}}{{\sum\nolimits_{o\in{\cal O}}{{e^{\sigma_{e}^{2}\left({s,o,{a_{o}}}\right)}}}}}\left|{{}_{\forall o\in{\cal O}}}\right.,\end{split}$ (23)
@@
10:     Get $s_{t+1}$ and $r^{i}_{t}$ from environment, for $i\in~{}[1,\cdots,N]$.
11:     Store $\{s_{t},(o_{t},a_{o,t}),\left[r^{1}_{t},\cdots,r^{N}_{t}\right],s_{t+1})\}$ into $D$.
@@
13:     Compute ${{\cal L}_{t}}(\theta_{ij}^{Q}),{{\cal L}_{t}}(\theta^{\mu})$, according to Eq. (14)(15).
14:     Update $\theta_{ij,t+1}^{Q}\leftarrow\theta_{ij,t}^{Q}-\alpha_{t}{\nabla}{{\cal L}_{t}}(\theta_{ij}^{Q})$.
15:     Update $\theta_{ij,t+1}^{\mu}\leftarrow\theta_{ij,t}^{\mu}-\beta_{t}{\nabla}{{\cal L}_{t}}(\theta^{\mu})$.
16:     $\theta_{ij,t+1}^{Q^{\prime}}\leftarrow{\tau}{\theta_{ij,t}^{Q}}+(1-{\tau}){\theta_{ij,t}^{Q^{\prime}}}$, $\theta_{t+1}^{\mu^{\prime}}\leftarrow{\tau}{\theta_{t}^{\mu}}+(1-{\tau}){\theta_{t}^{\mu^{\prime}}}$.
17:     update $\varsigma$, $s_{t}\leftarrow s_{t+1}$.
18:     if $s_{t}$ is terminal then
19:         Reset environment.
20:     end if
21:end for
22:return
@@ IV Implementation
This section presents the implementation details of using HPA-MoEC to perform the AD lane-changing task, including MDP formulation, training setup, and comparison models.
@@ IV-A1 State Space
For the agent to learn a driving policy that can handle the lane-changing task, the state space design must include all relevant factors that could impact lane-changing behavior. Specifically, the state space includes feature information about the Ego Vehicle (EV) and six surrounding vehicles (SVs) in its current and adjacent lanes:
@@ IV-A2 Hybrid Parameterized Action Space
For the lane-changing task, we design explicit hybrid parameterized actions as follows: i) discrete semantic decision action $b$, ii) continuous parameter $l$ for constructing a guiding path, and iii) continuous acceleration command $acc$. The concrete correspondence is: $b\leftarrow o,(l,acc)\leftarrow a_{o}$. Specifically, $b$ is selected from a discrete set $\left\{{LLC:-{w_{r}},RLC:{w_{r}},LK:0}\right\}$, where $w_{r}$ represents the road width, with $LLC$ and $RLC$ representing left and right lane-change, respectively, and $LK$ indicating lane-keeping. Considering the vehicle kinematic model [31], the value range for $l$ is defined as follows:
@@
where $R_{0}$ and $acc_{\text{max}}^{-}$ represent the minimum turning radius and maximum braking acceleration of the EV. In addition, the range of the acceleration command $acc$ is $\left[{-3{\rm{m/}}{{\rm{s}}^{2}},3{\rm{m/}}{{\rm{s}}^{2}}}\right]$.
@@
where $(x_{t+h},y_{t+h})$ represents the position of the point at time step $t+h$, and $H_{p}$ is the planning horizon. The coefficients $\gamma_{m}$ of the polynomial curve can be obtained by solving a system of linear equations. Specifically, the EV’s position and heading at the starting point are known, while the heading at the end point can be obtained from the road information [6]. Actually, the guiding path is determined by the selection of its endpoint position $(x_{t+H_{p}},y_{t+H_{p}})$, which is derived from the RL agent’s output, where $x_{t+H_{p}}=l_{t}$ and $y_{t+H_{p}}=b_{t}$. As the guiding path generated, the EV’s steering angle command $\delta$ is output using prior knowledge, specifically the Stanley algorithm in this paper. Finally, both the steering angle $\delta$ and the acceleration $acc$ are used together for EV driving control.
@@ IV-B Training Setup
We developed a three-lane structured road environment using the AD simulation platform, highway-env [51], in which the EV attempts to accomplish a multi-objective compatible lane-changing driving task. Specifically, all vehicles, including the EV, are randomly placed on the three-lane road with random initial speeds. The IDM and MOBIL models are applied to control the longitudinal and lateral movements of the SVs [31]. The SVs may change lanes at appropriate times to get closer to the target speed, potentially disrupting the EV. Additionally, we use the vehicle capacity (V/C) to represent traffic congestion, setting it to 0.5 to create moderate congestion. This ensures the EV has enough space to change lanes without oversimplifying the environment.
@@
Additionally, our method is tested on 200 episodes in both the training environment and the HighD [52] real-world dataset. For testing on the HighD dataset, the trained agent controls the vehicles exhibiting lane-changing behavior, while the SVs follow their predefined trajectories.
@@ IV-C1 Comparison Baseline
To comprehensively evaluate the proposed HPA-MoEC, we compare it with several widely used RL methods for the AD lane-changing task. To ensure fairness, all methods share the same training and testing environments, as well as the state space. The main difference is that, unlike HPA-MoEC, the other methods couple the attributes into a single reward function: ${{\cal R}_{base}}={\omega_{1}}{{\cal R}_{safe}}+{\omega_{2}}{{\cal R}_{gen}}$. More importantly, the action spaces structure and policy exploration strategies in the following methods differ:
@@ IV-C2 Ablation Model
To further validate the effectiveness of the three key techniques used in HPA-MoEC: i) hybrid parameterized actions with finer-grained abstract guidance; ii) a multi-objective compatible policy evaluation architecture; and iii) epistemic uncertainty-based policy exploration, we design the following ablation baselines:
@@
  • HPA: By further removing component ii, only one overall objective remains, with one corresponding critic for evaluating policies that considers multiple attributes.
  • HPA w/o GP: By further removing component i, this baseline generates only coarse-grained discrete semantic decisions as abstract guidance, which are combined with the PID controller to output steering angles. It retains a hybrid action space that also provides continuous acceleration commands.
@@ V-B1 Testing with Rule-Based SVs
For safety performance, HPA-MoEC demonstrates the lowest CR, second only to SAC-C, highlighting its strong focus on safety. This is facilitated by a policy evaluation design with safety attribute as a separate objective, achieving a CR reduction of 67% and 69% for HPA compared to SAC-H and PPO-H, respectively. Notably, SAC-C adopts a highly conservative driving policy, greatly reducing the CR at the cost of driving efficiency. Additionally, DQN has a CR of 0.38%, much higher than other methods. With an average of 7.73 NL per episode, this indicates that its more aggressive driving policy increases the risk of putting the EV in danger.
@@ V-C1 Training Performance
For HPA w/o GP, there are even lower rewards and higher CR compared to HPAs. This indicates that the introduction of finer-grained guiding paths enhances the connection between the agent’s outputs and driving behavior, leading to further improvements in both general policy performance and safety.
@@ V-C2 Testing with Rule-Based SVs
HPA w/o GP performs the worst across all metrics compared to the other ablation baselines. Notably, removing the guiding path results in approximately a 100% increase in VS compared to HPA, highlighting the larger fluctuations in lateral driving behavior. In addition, its AS decreases by 14%, with a wider distribution, while the CR increases by 38%, reflecting a decline in both driving efficiency and safety. Therefore, implementing a hybrid parameterized action space with finer-grained guidance paths helps the agent promote multi-objective driving, particularly in terms of reducing fluctuations in driving behavior.
@@ V-C3 Testing in HighD-Dataset
HPA w/o GP 0.946 14.43 6.08 0.0041 0.077 0.06%
The testing results for all ablation baselines in the HighD dataset are shown in Fig. 8 and Table V. HPA-Mo falls slightly below HPA-MoEC in driving efficiency, but both have good driving performance. In contrast, HPA lags clearly behind both previous methods in AS and NL and has a higher CR. The ’HPA w/o GP’ is even worse, accompanying a notable increase in VS. This suggests that the multi-objective policy evaluation architecture and the hybrid parameterized action space with guiding paths still promote the compatibility of the objectives of driving efficiency, action consistency and safety in the HighD dataset.
@@ V-D Discussion
In summary, our HPA-MoEC method outperforms all the RL comparison baselines, where all three key technology components play a significant role in facilitating the learning of a multi-objective compatible policy. The hybrid parameterized action enhances the connection between agent actions and driving behavior by simultaneously outputting finer-grained guiding paths as well as direct acceleration commands. This action space structure promotes multi-objective compatibility, particularly enhancing action consistency by reducing driving behavior fluctuations while maintaining flexibility. The multi-objective policy evaluation architecture guides the agent in improving policy learning by treating general and safety attributes as distinct objectives and building the corresponding reward function and critic. This policy evaluation architecture improves both the driving general performance and safety, demonstrating its ability to achieve multi-objective compatible driving. In addition, the epistemic uncertainty-based policy exploration mechanism accelerates the convergence of multi-objective compatible viable policies, improving the training efficiency. Furthermore, the testing in HighD demonstrates that HPA-MoEC remains superior under real scenario datasets and possesses better application prospects.
@@
Additionally, to better observe the impact of our exploration mechanism on epistemic uncertainty, we denote ’w/o $-\cal E$ as an attempt. In this attempt, ensemble-critics generate epistemic uncertainty but do not use it for exploration, instead performing random exploration. The curves in Figure 1 show how epistemic uncertainty evolves throughout the policy improvement process. Our HPA-MoEC experiences higher average uncertainty in the early training phases, and then makes the uncertainty lower more rapidly during exploration. This suggests that HPA-MoEC explores more fully while converging the policy faster than randomized exploration. Further, the changes in epistemic uncertainty for the three lane-change decisions follow a similar trend. Notably, changing lanes—whether to the left or right—results in higher uncertainty compared to lane keeping, suggesting that lane changes involve greater unknowns and risks.
@@ VI Conclusion and Future Work
This paper proposes a Multi-objective Ensemble-Critic (HPA-MoEC) reinforcement learning method with Hybrid Parameterized Action space, capable of efficiently learning multi-objective compatible driving policies. Our method includes three key components: i) the hybrid parameterized action space that simultaneously generates abstract guidance and concrete control commands, ii) the multi-objective compatible policy evaluation framework that considers multiple driving attributes, and iii) the epistemic uncertainty-based policy exploration strategy. We conduct the training and testing of the policy in both simulated traffic environments and the HighD dataset. The results show that HPA-MoEC effectively learns a multi-objective compatible autonomous driving policy in terms of efficiency, action consistency, and safety. The ablation study further demonstrated the role of the three technology components in HPA-MoEC in promoting multi-objective compatibility.
Future work aims to further investigate how uncertainty can be leveraged in autonomous driving, particularly to encourage more cautious behavior during testing.

### v3

@@ III-C Uncertainty Estimation and Exploration Strategy
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
10:  Get $s_{t+1}$ and $r_{i,t}$ from environment, for $i\!\in\![1,\!\cdots\!,\!N]$.
11:  Store $\{s_{t},(o_{t},a_{o,t}),\left[r_{1,t},\cdots,r_{N,t}\right],s_{t+1})\}$ into $D$.
@@
13:  Calculate ${{\cal L}_{t}}(\theta_{ij}^{Q})$ for each critic, according to Eq.(14).
14:  Update every critic network via gradient descent:
15:      $\theta_{ij,t+1}^{Q}\leftarrow\theta_{ij,t}^{Q}-\alpha_{t}{\nabla}{{\cal L}_{t}}(\theta_{ij}^{Q})$.
16:  Calculate ${{\cal L}_{t}}(\theta^{\mu})$ for actor, according to Eq.(15).
17:  Update actor network via gradient descent:
18:      $\theta_{ij,t+1}^{\mu}\leftarrow\theta_{ij,t}^{\mu}-\beta_{t}{\nabla}{{\cal L}_{t}}(\theta^{\mu})$.
19:  Soft-update every target critic network:
20:      $\theta_{ij,t+1}^{Q^{\prime}}\leftarrow{\tau}{\theta_{ij,t}^{Q}}+(1-{\tau}){\theta_{ij,t}^{Q^{\prime}}}$.
21:  Soft-update actor:
22:      $\theta_{t+1}^{\mu^{\prime}}\leftarrow{\tau}{\theta_{t}^{\mu}}+(1-{\tau}){\theta_{t}^{\mu^{\prime}}}$.
23:  update $\varsigma$, $s_{t}\leftarrow s_{t+1}$.
24:  if $s_{t}$ is terminal then
25:   Reset environment.
26:  end if
27:end for
28:return
@@ IV Implementation
Multi-lane highway scenarios are both common and challenging, requiring driving policies that satisfy objectives such as efficiency, action consistency, and safety. This section presents the implementation details of HPA-MoEC in these scenarios, including the MDP formulation, training setup, and baseline models.
@@ IV-A1 State Space
An appropriate state space representation is essential for effective policy learning. Specifically, the state space includes feature information about the Ego Vehicle (EV) and six surrounding vehicles (SVs) in its current and adjacent lanes:
@@ IV-A2 Hybrid Parameterized Action Space
For multi-lane scenarios with hybrid road modalities, we design explicit hybrid parameterized actions as follows: i) discrete semantic decision action $b$, ii) continuous parameter $l$ for constructing a guiding path, and iii) continuous acceleration command accacc. The concrete correspondence is: $b\leftarrow o,(l,acc)\leftarrow a_{o}$. Specifically, $b$ is selected from a discrete set $\left\{{LLC:-{w_{r}},RLC:{w_{r}},LK:0}\right\}$, where $w_{r}$ represents the road width, with $LLC$ and $RLC$ representing left and right lane-change, respectively, and $LK$ indicating lane-keeping. Considering the vehicle kinematic model [16], the value range for $l$ is defined as follows:
@@
where $R_{0}$ and $acc_{\text{max}}^{-}$ represent the minimum turning radius and maximum braking acceleration of the EV. In addition, the range of the acceleration command accacc is $\left[{-3{\rm{m/}}{{\rm{s}}^{2}},3{\rm{m/}}{{\rm{s}}^{2}}}\right]$.
@@
where $(x_{t+h},y_{t+h})$ represents the position of the point at time step $t+h$, and $H_{p}$ is the planning horizon. The coefficients $\gamma_{m}$ of the polynomial curve can be obtained by solving a system of linear equations. Specifically, the EV’s position and heading at the starting point are known, while the heading at the end point can be obtained from the road information [6]. Actually, the guiding path is determined by the selection of its endpoint position $(x_{t+H_{p}},y_{t+H_{p}})$, which is derived from the RL agent’s output, where $x_{t+H_{p}}=l_{t}$ and $y_{t+H_{p}}=b_{t}$. As the guiding path generated, the EV’s steering angle command $\delta$ is output using prior knowledge, specifically the Stanley algorithm in this paper. Finally, both the steering angle $\delta$ and the acceleration accacc are used together for EV driving control.
@@ IV-B Training Setup
Figure 3: Schematic diagram of multi-lane highway environments in highway-env. Green vehicles represent EVs, while blue vehicles represent SVs.
We developed a three-lane structured road environment using the AD simulation platform, highway-env [56], in which the EV attempts to accomplish a multi-objective compatible driving task. A schematic diagram of the study scenario is shown in Fig. 3. Specifically, all vehicles, including the EV, are randomly placed on the three-lane road with random initial speeds. The IDM and MOBIL models are applied to control the longitudinal and lateral movements of the SVs [16]. The SVs may change lanes at appropriate times to get closer to the target speed, potentially disrupting the EV. Additionally, we use the vehicle capacity (V/C) to represent traffic congestion, setting it to 0.5 to create moderate congestion. This ensures the EV has enough space to change lanes without oversimplifying the environment.
@@
Additionally, our method is tested on 200 episodes in both the training environment and the HighD [57] real-world dataset. For testing on the HighD dataset, the trained agent controls randomly selected vehicles, while the SVs follow their predefined trajectories.
@@ IV-C1 Comparison Baseline
To comprehensively evaluate the proposed HPA-MoEC, we compare it with several widely used RL methods for the AD task. All methods share the same training and testing environments, as well as the state space. The main difference is that, unlike HPA-MoEC, the other methods couple the attributes into a single reward function: ${{\cal R}_{base}}={\omega_{1}}{{\cal R}_{safe}}+{\omega_{2}}{{\cal R}_{gen}}$. More importantly, the action spaces structure and policy exploration strategies in the following methods differ:
@@
To ensure fair comparisons and reliable conclusions, all methods use the same network architecture, learning rate, and other key hyperparameters. Additionally, for method-specific parameters, we perform extensive tuning within reasonable ranges and select the optimal configuration for each method.
@@ IV-C2 Ablation Model
To further validate the effectiveness of the three key techniques used in HPA-MoEC: i) Hybrid parametric action space structure; ii) Multi-critic policy evaluation architecture; and iii) Epistemic uncertainty-based policy exploration, we design the following ablation baselines:
@@
  • HPA: By removing component ii from HPA-Mo, only one overall objective remains, with one corresponding critic for evaluating policies that considers multiple attributes. In fact, this baseline is similar to the algorithm in [16].
  • DA-Mo: By removing component i from HPA-Mo, this baseline generates only coarse-grained discrete semantic decisions as abstract guidance, which are combined with the PID controller to output steering angles. In fact, this baseline is similar to part of the work in [21].
@@ V-B1 Testing with Rule-Based SVs
For safety performance, HPA-MoEC demonstrates the second-lowest CR, trailing only to SAC-C, highlighting its strong focus on safety. This is facilitated by a policy evaluation design with safety attribute as a separate objective, achieving a CR reduction of 67% and 69% for HPA compared to SAC-H and PPO-H, respectively. Notably, since SAC-C directly outputs control commands through a single continuous action space, it learns an over-conservative driving policy. Although this extreme concern for short-term safety significantly reduces CR, it greatly sacrifices efficiency and action consistency. In contrast, the slight compromise in safety offered by HPA-MoEC brings significant improvements in efficiency and action consistency, which is more aligned with the multi-objective requirements of AD. Additionally, DQN has a CR of 0.38%, much higher than other methods. With an average of 7.73 NL per episode, this indicates that its more aggressive driving policy increases the risk of putting the EV in danger.
@@ V-C1 Training Performance
For DA-Mo, the reward it can obtain when converging is the lowest, and the CR is the highest. This shows that hybrid action space structure plays an important role in improving policy execution capabilities. A finer-grained guidance path enhances the correlation between agent output and driving behavior, further improving overall policy performance and safety.
@@ V-C2 Testing with Rule-Based SVs
DA-Mo performs the worst across all metrics compared to the other ablation baselines. Notably, removing the hybrid action space results in approximately a 25% increase in VS compared to HPA, highlighting the larger fluctuations in lateral driving behavior. In addition, its AS decreases by 15%, with a wider distribution, while the CR increases by 100%, reflecting a decline in both driving efficiency and safety. Therefore, implementing a hybrid parameterized action space with finer-grained guidance paths helps the agent promote multi-objective driving, particularly in terms of reducing fluctuations in driving behavior.
@@ V-C3 Testing in HighD-Dataset
DA-Mo 0.947 14.68 6.53 0.0036 0.076 0.04%
The testing results for all ablation baselines in the HighD dataset are shown in Fig. 9 and Table V. HPA-Mo falls slightly below HPA-MoEC in driving efficiency, but both have good driving performance. In contrast, HPA lags clearly behind both previous methods in AS and NL and has a higher CR. The DA-Mo is even worse, accompanying a notable increase in VS. This suggests that the multi-objective policy evaluation architecture and the hybrid parameterized action space with guiding paths still promote the compatibility of the objectives of driving efficiency, action consistency and safety in the HighD dataset.
@@ V-D Discussion
In summary, our HPA-MoEC method outperforms all the RL comparison baselines, where all three key technology components play a significant role in facilitating the learning of a multi-objective compatible policy. The hybrid parameterized action enhances the connection between agent actions and driving behavior by simultaneously outputting finer-grained guiding paths as well as direct acceleration commands. This action space structure promotes multi-objective compatibility, particularly enhancing action consistency by reducing driving behavior fluctuations while maintaining flexibility. The multi-objective policy evaluation architecture guides the agent in improving policy learning by treating general and safety attributes as distinct objectives and building the corresponding reward function and critic. This policy evaluation architecture improves both the driving general performance and safety, demonstrating its ability to achieve multi-objective compatible driving. In addition, the epistemic uncertainty-based policy exploration mechanism accelerates the convergence of multi-objective compatible viable policies, improving the training efficiency. It is also noteworthy that SVs in the HighD dataset exhibit human driving behaviors, which differ significantly from those in simulation traffic. Although HPA-MoEC is trained in simulated traffic, it still achieves strong driving performance when confronted with unfamiliar SVs. This demonstrates that HPA-MoEC possesses strong generalization capabilities and can effectively adapt to unfamiliar environments.
@@
[…]
@@ VI Conclusion and Future Work
This paper proposes a Multi-objective Ensemble-Critic (HPA-MoEC) reinforcement learning method with Hybrid Parameterized Action space, capable of efficiently learning multi-objective compatible driving policies. HPA-MoEC adopts a more advanced MORL architecture, in which multiple reward functions guide ensemble-critics to focus on specific driving objectives. Meanwhile, the architecture integrates a hybrid parameterized action space structure, which can simultaneously generate abstract guidance and specific control commands that fit the hybrid road modality. In addition, an uncertainty-based exploration mechanism is developed to achieve faster learning of multi-objective compatible policies. We conduct the training and testing of the policy in both simulated traffic environments and the HighD dataset. The results show that HPA-MoEC effectively learns a multi-objective compatible autonomous driving policy in terms of efficiency, action consistency, and safety. The ablation study further demonstrated the role of technology components in HPA-MoEC in promoting multi-objective compatibility.
One limitation of our study is that the driving scenarios for training and testing are restricted to multi-lane highways. Although this typical structured road environment differs from other road types such as ramps and intersections, the driving objectives of EV in these various scenarios are generally similar: selecting appropriate behavioral goals and interacting with other vehicles. The key difference lies in how the state space is designed to enable the RL agent to comprehensively perceive the environment. Therefore, in future work, we aim to use higher-dimensional perception information (such as BEV images) as the state space to extend the application of HPA-MoEC to more complex traffic scenarios.